solve X^4-23x^2-108=0




thanks

Answer by mananth(16949)

x^4-23x^2-108=0
...
x^4-27x^2+4x^2-108=0
x^2(x^2-27)+4(x^-27)=0
(x^2-27)(x^2+4)=0
...
m.ananth@hotmail.ca